OpenStudy (anonymous):

3x + 4 = 19 , { 4,5,6,7} A. 4 B. 5 C. 6 D. 7

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (freckles):

plug in the numbers from the set and see which gives you a true equation

OpenStudy (freckles):

for example is 3*4+4=19 true?

OpenStudy (anonymous):

isnt it 16?...

OpenStudy (freckles):

so is 16=19 true or false?

OpenStudy (freckles):

is 16 the same number as 19?

OpenStudy (anonymous):

False

OpenStudy (freckles):

right plug in the next number in the set

OpenStudy (anonymous):

i think its 7

OpenStudy (freckles):

is 3*5+4=19 a true equation

OpenStudy (freckles):

so you think 3*7+4=19 is true ?

OpenStudy (anonymous):

no

OpenStudy (freckles):

you have 3x+4=19 you are just trying to see which of the possible x values 4,5,6, or 7 are going to give you a true equation here: 3*4+4=19 is false 3*5+4=19 is ? 3*6+4=19 is ? 3*7+4=19 is false

OpenStudy (anonymous):

3*5+4 ?

OpenStudy (anonymous):

5?

OpenStudy (freckles):

3*5 is 15 and 3*5+4=15+4 which is 19 19=19 is totally true because 19 is 19

OpenStudy (anonymous):

so its that one the answer?

OpenStudy (freckles):

did you read this from above: "you are just trying to see which of the possible x values 4,5,6, or 7 are going to give you a true equation here:"

OpenStudy (freckles):

19=19 is what plugging in 5 gave us we are already determine 19=19 is true what doubts do you still have?

OpenStudy (anonymous):

ok thanks...

OpenStudy (freckles):

np :) so you are totally sure the answer is...

OpenStudy (anonymous):

5

OpenStudy (freckles):

right! :)

OpenStudy (trojanpoem):

3x + 4 = 19 let 4 be in the other side with negative sign 3x = 19 - 4 = 15 3x = 15 divide over 3 x = 5

OpenStudy (anonymous):

thanks!!